OpenStudy (anonymous):

Please Help! Determine which ratio forms a proportion with 10/6 by using cross products

OpenStudy (anonymous):

what are the given ratios?

OpenStudy (anonymous):

15/12 and 12/30 and 20/9 and 30/18

jimthompson5910 (jim_thompson5910):

Let's assume that 10/6 = 15/12 is a true equation. If it were, then we can cross multiply to get 10/6 = 15/12 10*12 = 6*15 120 = 90 since the last equation is false, this means that the original equation is false.

jimthompson5910 (jim_thompson5910):

You can use that to determine if any two fractions are equal

OpenStudy (anonymous):

\[\frac{ 10 }{ 6 } = \frac{ x }{ y }\] \[10 *y = 6*x\]

OpenStudy (anonymous):

for any x find the corresponding value of y. There are millions of values.